# Learning-based agricultural management in partially observable environments subject to climate variability

### v1

@@ 3.2 Results and discussions
Year
MDP-28
POMDP-28
MDP-10
POMDP-10
1965
235
350
187
350
1980
594
612
460
612
1999
515
584
435
584
2020
1435
1471
1200
1466
@@
Policy from
Yield (kg/ha)
Nitrogen input (kg/ha)
Nitrate leaching (kg/ha)
MDP-28
9247
360
0.14
POMDP-28
9243
180
0.12
MDP-10
9226
560
0.20
POMDP-10
9243
180
0.12
Expert policy 1
6236
56
0.12
Expert policy 2
9247
224
0.26
@@
In accordance with a 1999 report titled ’Iowa Crop Performance Test- Corn District 2,’ published by Iowa State University, corn yielded for various brands ranged between 146 bu/acre and 192 bu/acre, with an average yield of 169.4 bu/acre. It is worth noting that these numbers represented yields of wet corn. To make accurate comparisons with our simulation results, these wet yields must be converted to dry corn yields. The report specified an average moisture content of 17.5% for the wet corn. In our conversion process, we use a moisture content of 14% for dry corn, as referenced in a study by Yakoub et al. Yakoub2017 . After converting to dry yield, the range of corn yield in 1999 spans from 8507.53 kg/ha to 11197.74 kg/ha, with an average yield of 9868.59 kg/ha.
Our simulation results, averaging approximately 9243 kg/ha from optimal policies in Table 4, exhibit a deviation of about 6.4% from the actual yield. This aligns with variations observed in some prior studies (e.g., 13% and 9.5%) that compared the DSSAT simulation results with the actual crop productions Yakoub2017 ; Nouna2000 . It is worth noting that the DSSAT maize models, specifically CERES, were initially published 37 years ago Jones1986 . As such, they may not perfectly capture modern maize production methods, which have evolved substantially over the years. Nonetheless, the optimal policies derived from POMDP models recommend reduced nitrogen usage (as shown in Table 4) compared to the second expert policy, which we will use as a baseline for our subsequent studies in this paper.
In order to assess the influence of nitrate leaching on both nitrogen fertilization practices and corn yield, we conduct a series of simulations to obtain optimal policies from the POMDP-10 model involving different values of $w_{3}$, ranging from 0 to 50 times the value of $w_{2}$. Throughout the comparisons, the overall quantity of nitrogen usage remains constant, while the timing of fertilization applications is adjusted according to the specified $w_{3}$ values. Upon closer examination, we find that assigning a higher weight to nitrate leaching in the reward function (as defined in Equation (6)) leads to a reduction in nitrate leaching. This reduction is achieved by minimizing fertilization on rainy days whenever possible. For the subsequent studies detailed in this paper, we choose to maintain $w_{3}$ five times of $w_{2}$ in the reward function. This decision aligns with the approach used in a previously referenced study Wu2022 .
@@ 4.1 Impact of higher temperature
Over the past 70 years, the global climate patterns have undergone significant changes primarily driven by anthropogenic activities, with temperature increases being one of the most prominent aspects Rabatel2013 . Historical data reveals a steady rise in average temperatures, within a 0.98-degree Celsius increase since 1880. Notably, this warming trend has accelerated in recent decades, with a 0.94-degree Celsius rise recorded in the last 60 years alone. In this study, we systematically elevated the daily average temperature from the 1999 baseline by increments of 0.5, 1, 2, and 5 degrees Celsius NASA throughout the year while following a consistent pattern. The precipitation remains the same as in 1999, so we can investigate the impact of temperature variation on fertilization management and corn yield.
@@
Nitrate leaching (kg/ha) 0.11 0.10 0.28
$+1^{\circ}$
@@
Nitrate leaching (kg/ha) 0.10 0.10 0.29
$+2^{\circ}$
@@
Nitrate leaching (kg/ha) 0.09 0.09 0.29
$+5^{\circ}$
@@
Our simulations run from April 10th to October 30th, with an average air temperature of 10.9 degrees Celsius in 1999 on the planting day (May 27th), aligning with the recommended corn planting temperature of above 10 degrees Celsius by experts Abendroth2017 . However, when temperatures surge by more than 2.5 degrees, corn yield starts to decline, possibly indicating that the heat becomes too intense for healthy corn growth. This hypothesis gains further support when the temperature is raised by 5 degrees Celsius, resulting in a significant drop in corn yield to just half its original volume under the actual weather conditions of 1999.
While examining the data presented in Table 5, it becomes evident that various policies exhibit similar trends in corn yield as daily temperature systematically increases. However, notable differences emerge in nitrogen input when corn yield starts to decline significantly due to substantial temperature escalation. The ‘optimal policies,’ which consistently employ significantly less nitrogen than the ‘1999 policy,’ lead to higher rewards. Specifically, when the temperature increases by 5 degrees, the ’optimal policy’ yields a reward (representing the net income) 30% higher than the ‘1999 policy.’ In contrast, the expert policy maintains the same fertilizer usage but results in substantially higher nitrate leaching compared to other policies.
@@ 4.2 Impact of insufficient precipitation
Table 6 presents the simulated outcomes under varying precipitation scenarios. Similar to Table 5, this table compares corn yield, nitrogen input, and nitrate leaching between various policies, including the ‘1999 policy,’ optimal policies, and an expert policy. It can be seen that when precipitation decreases, the corn yield is significantly impacted. The corn yield can be less than half the standard value when the weather is severely dry. Given that corn is a moisture-intensive crop Paolo2008 , the results are convincing.
@@ 4.3 Heat wave and drought
To further study the impact of extreme weather events on agriculture and fertilization management, we consider two real scenarios that occurred in Iowa: the heat wave in 1983 and the drought in 1988. Previous research Motha2005 indicated that these extreme weather events led to 32% and 38.5% reduction in Iowa’s corn yields compared to the previous years.
@@
Figures  4 and 5 depict the comparisons of monthly average temperatures and precipitations for the months of April through October in 1982 and 1983. It can be seen that the average temperatures in June, July, August, and September of 1983 were higher than in 1982 by 3.1 degrees, 1.7 degrees, 4.3 degrees, and 1.5 degrees, respectively. However, the overall precipitation in 1983 was higher than in 1982, especially in June 1983. Therefore, there was a heatwave in 1983 but not a drought.
@@ 5 Conclusion and future works
CRediT authorship contribution statement
Zhaoan Wang: Conceptualization, Data curation, Investigation, Methodology, Writing – original draft. Shaoping Xiao: Conceptualization, Methodology, Supervision, Funding acquisition, Writing - review & editing. Junchao Li: Methodology, Writing - review & editing. Jun Wang: Conceptualization, Funding acquisition, Writing - review & editing.
@@ Acknowledgments
The authors received support from the University of Iowa OVPR Interdisciplinary Scholars Program for this study. Z.W., J.L., and S.X. received supports from the U.S. Department of Education (E.D. #P116S210005) and the National Science Foundation (#2226936).

### v2

@@ 3.2 Results and discussions
Year MDP-28 POMDP-28 MDP-10 POMDP-10
1965 235 350 187 350
1980 594 612 460 612
1999 515 584 435 584
2020 1435 1471 1200 1466
@@
Policy from Yield (kg/ha) Nitrogen input (kg/ha) Nitrate leaching (kg/ha)
MDP-28 9247 360 0.14
POMDP-28 9243 180 0.12
MDP-10 9226 560 0.20
POMDP-10 9243 180 0.12
Expert policy 1 6236 56 0.12
Expert policy 2 9247 224 0.26
@@
In accordance with a 1999 report titled ’Iowa Crop Performance Test- Corn District 2,’ published by Iowa State University, corn yielded for various brands ranged between 146 bu/acre and 192 bu/acre, with an average yield of 169.4 bu/acre. It is worth noting that these numbers represented yields of wet corn. To make accurate comparisons with our simulation results, these wet yields must be converted to dry corn yields. The report specified an average moisture content of 17.5% for the wet corn. In our conversion process, we use a moisture content of 14% for dry corn, as referenced in a study by Yakoub et al. [35]. After converting to dry yield, the range of corn yield in 1999 spans from 8507.53 kg/ha to 11197.74 kg/ha, with an average yield of 9868.59 kg/ha.
Our simulation results, averaging approximately 9243 kg/ha from optimal policies in Table 4, exhibit a deviation of about 6.4% from the actual yield. This aligns with variations observed in some prior studies (e.g., 13% and 9.5%) that compared the DSSAT simulation results with the actual crop productions [35, 36]. It is worth noting that the DSSAT maize models, specifically CERES, were initially published 37 years ago [37]. As such, they may not perfectly capture modern maize production methods, which have evolved substantially over the years. Nonetheless, the optimal policies derived from POMDP models recommend reduced nitrogen usage (as shown in Table 4) compared to the second expert policy, which we will use as a baseline for our subsequent studies in this paper.
In order to assess the influence of nitrate leaching on both nitrogen fertilization practices and corn yield, we conduct a series of simulations to obtain optimal policies from the POMDP-10 model involving different values of $w_{3}$, ranging from 0 to 50 times the value of $w_{2}$. Throughout the comparisons, the overall quantity of nitrogen usage remains constant, while the timing of fertilization applications is adjusted according to the specified $w_{3}$ values. Upon closer examination, we find that assigning a higher weight to nitrate leaching in the reward function (as defined in Equation (6)) leads to a reduction in nitrate leaching. This reduction is achieved by minimizing fertilization on rainy days whenever possible. For the subsequent studies detailed in this paper, we choose to maintain $w_{3}$ five times of $w_{2}$ in the reward function. This decision aligns with the approach used in a previously referenced study [17].
@@ 4.1 Impact of higher temperature
Over the past 70 years, the global climate patterns have undergone significant changes primarily driven by anthropogenic activities, with temperature increases being one of the most prominent aspects [38]. Historical data reveal a steady rise in average temperatures, with a 0.98-degree Celsius increase since 1880. Notably, this warming trend has accelerated in recent decades, with a 0.94-degree Celsius rise recorded in the last 60 years alone. In this study, we systematically elevated the daily average temperature from the 1999 baseline by increments of 0.5, 1, 2, and 5 degrees Celsius [39] throughout the year while following a consistent pattern. The precipitation remains the same as in 1999, so we can investigate the impact of temperature variation on fertilization management and corn yield.
@@
Nitrate leaching (kg/ha) 0.11 0.10 0.18
$1^{\circ}$C
@@
Nitrate leaching (kg/ha) 0.10 0.10 0.12
$2^{\circ}$C
@@
Nitrate leaching (kg/ha) 0.09 0.09 0.09
$5^{\circ}$C
@@
Our simulations run from April 10th to October 30th, with an average air temperature of 10.9 degrees Celsius in 1999 on the planting day (May 27th), aligning with the recommended corn planting temperature of above 10 degrees Celsius by experts [40]. However, when temperatures surge by more than 2.5 degrees, corn yield starts to decline, possibly indicating that the heat becomes too intense for healthy corn growth. This hypothesis gains further support when the temperature is raised by 5 degrees Celsius, resulting in a significant drop in corn yield to just half its original volume under the actual weather conditions of 1999.
While examining the data presented in Table 5, it becomes evident that various policies exhibit similar trends in corn yield as daily temperature systematically increases. However, notable differences emerge in nitrogen input when corn yield starts to decline significantly due to substantial temperature escalation. The ‘optimal policies,’ which consistently employ significantly less nitrogen than the ‘1999 policy,’ lead to higher rewards. Specifically, when the temperature increases by 5 degrees, the ’optimal policy’ yields a reward (representing the net income) 30% higher than the ‘1999 policy.’
@@ 4.2 Impact of insufficient precipitation
Table 6 presents the simulated outcomes under varying precipitation scenarios. Similar to Table 5, this table compares corn yield, nitrogen input, and nitrate leaching between various policies, including the ‘1999 policy,’ optimal policies, and an expert policy. It can be seen that when precipitation decreases, the corn yield is significantly impacted. The corn yield can be less than half the standard value when the weather is severely dry. Given that corn is a moisture-intensive crop [41], the results are convincing.
@@
The notably low nitrate leaching values of 0.001 kg/ha observed under both severe precipitation reductions may reflect not only the physical constraints of water-limited nitrogen movement in soils but also a limitation of the DSSAT simulator under such extreme environmental changes. DSSAT’s leaching algorithms are primarily parameterized for conditions within a plausible range of soil moisture and temperature dynamics derived from historical and experimental datasets. When climate conditions deviate far beyond these calibration bounds the model may underestimate nitrate transport because simulated soil water fluxes become effectively negligible. In reality, even under such stressed conditions, small but measurable nitrate losses can occur through occasional rainfall events, preferential flow in cracked soils, or subsurface lateral movement. These pathways may not be fully captured by DSSAT’s standard routines without site-specific calibration. For this reason, we retain the model’s output of 0.001 kg/ha as the reported value, acknowledging it as the simulator’s minimal positive output rather than a definitive measurement of field leaching.
@@ 4.3 Heat wave and drought
To further study the impact of extreme weather events on agriculture and fertilization management, we consider two real scenarios that occurred in Iowa: the heat wave in 1983 and the drought in 1988. Previous research [23] indicated that these extreme weather events led to 32% and 38.5% reduction in Iowa’s corn yields compared to the previous years.
@@
Figures 4 and 5 depict the comparisons of monthly average temperatures and precipitations for the months of April through October in 1982 and 1983. It can be seen that the average temperatures in June, July, August, and September of 1983 were higher than in 1982 by 3.1 degrees, 1.7 degrees, 4.3 degrees, and 1.5 degrees, respectively. However, the overall precipitation in 1983 was higher than in 1982, especially in June 1983. Therefore, there was a heatwave in 1983 but not a drought.
@@ 5 Conclusion and future works
Author contributions statement
Zhaoan Wang. Shaoping Xiao. and Jun Wang. Conceptualization, Zhaoan Wang. Data curation, Zhaoan Wang. Investigation, Zhaoan Wang. Shaoping Xiao. and Junchao Li. Methodology, Zhaoan Wang. Writing – original draft, Shaoping Xiao. Supervision, Shaoping Xiao. and Jun Wang. Funding acquisition, Shaoping Xiao. Junchao Li. and Jun Wang. Writing - review & editing.
All authors reviewed the manuscript.
@@
Funding
This material is based upon work supported by the U.S. Department of Education under Grant Number ED#P116S210005 and the National Science Foundation under Grant Numbers 2226936 and 2420405. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U.S. Department of Education and the National Science Foundation.
@@ Acknowledgments
The authors acknowledge the support from the University of Iowa OVPR Interdisciplinary Scholars Program for this study.